Write the equation of the line passing through each of the given pairs of points. Write your result in slope-intercept form, where possible. (-1,3) and (4,-2)
Answer by scott8148(6628)

slope is (-2-3)/(4-(-1)) or -5/5 or -1

using point-slope ... y-3=-1(x-(-1)) ... y-3=-x-1

in slope-intercept ... y=-x+2